5x - 4(x +9) = 2x - 3
Found 2 solutions by Jeannie1012, kaketto:
Answer by Jeannie1012(8)

To solve this it is important to remember order of operations.
5x-4(x+9) = 2x-3
Step 1: Distribute the multiplication in the middle -4(x+9)to get
5x-4x-36 = 2x-3
Step 2: Combine like terms where you can to get
x-36 = 2x-3
Step 3: Now we want all the x's on one side. I like to keep it positive when I can so we are going to subtract the x from both sides to get
-36 = x - 3
Step 4: We want the x alone on one side so now we need to add 3 to both sides to get that done and we end up with
-33 = x or x = -33

You also could have done this another way.
1) distribute to get 5x-4x-36 = 2x-3
2) combine like terms x-36 = 2x -3
3) instead of subtracting the x like we did above, subtract the 2x from both sides and get -x - 36 = -3
4) now add 36 to each side to get the x alone and you have -x = 33
if you went this way, which is fine you now need to divide by -1 so that x is positive and -x divided by -1 is x and 33 divided by -1 is -33
either way x = -33
check by plugging -33 in where x is in the equation 5x-4(x+9) = 2x -3 becomes
5(-33) -4(-33 +9) = 2(-33) -3 don't forget distribute the middle here as well.
-165 +132 - 36 = -66 - 3
-69 = -69
since that statement is true the answer checks.
Hope this helps. Good Luck!

Answer by kaketto(2)

5x - 4(x +9) = 2x - 3
The first step is to distribute the () part with FOIL (Front Outside Inside Last)
so that would be -4(x+9) keeping the negative factor of 4 and distributing this
that would be -4x + -36
then plug that back into the equation 5x-4x-36 = 2x-3
at this point we need to eliminate x on one side to solve for x
so first add like terms 5x-4x becomes x so we are left with x-36=2x-3
now get x on one side so let us subtract x from the left and also add 3 to the right side to get numbers on one side and x on other
x-36=2x-3
-x +3 = -x +3
so we are left with -33 = x as the answer